# thanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ I Introduction
The high-efficiency IFM is employed in various applications, such as quantum interrogation [6], quantum tomography [7, 8, 9], quantum computation [10, 11, 12], and quantum cryptography [13, 14]. Importantly, this physical phenomenon is highly suitable for secure communication known as counterfactual communication, wherein information is conveyed without the transmission of any particles through the channel. As shown in Fig. 1(a), in classical communication, the communicating parties (Alice and Bob) must transmit information via physical particles passing through the channel, which may be eavesdropped by Eve. In counterfactual communication, however, Alice emits a quantum state to herself, and Bob conveys information by altering the evolution outcome of this state via $\hat{P}$. Throughout the entire process, no particles pass through the transmission channel. Thus the eavesdropper Eve cannot intercept any particles from the channel, and consequently cannot obtain the information. Furthermore, by utilizing the quantum Zeno effect, the success rate of counterfactual communication can be increased to 100% in theory [15].
@@ II.1 Counterfactual communication protocol
Figure 1: The schematic and structure of the upgraded counterfactual communication protocol. (a) During the classical communication process, Eve can obtain the information from Alice to Bob by intercepting the information-carried particles. However, since Alice and Bob do not transmit any particles during the counterfactual communication, Eve can not acquire any information. (b) The updated protocol for the counterfactual communication. (c) The success probability of transmitting bit 0 and bit 1 are simulated with various $M$ and $N$. Balancing the trade-off between feasibility of implementation and success probabilities of transmission, we choose $M=3$ and $N=6$ in our experimental scheme.
@@
Therefore, the specific optical scheme and the operation of the upgraded protocol are as follows, shown in Fig. 1b. The outer cycles of interferometers consist of $M$ beam splitters ($BS_{M}$s) with the reflectivity of $cos^{2}(\pi/2M)$ each. Serving as one arm of the outer interferometer, the inner cycles contain two sets of tandem interferometers connected by a two-sided mirror. Each set of inner interferometers comprises $N$ beam splitters ($BS_{N}$s) with a reflectivity of $cos^{2}(\pi/2N)$ for each $BS_{N}$. In accordance with the quantum Zeno effect, the photon state remains stable within the nested cycles of interferometers. The structural unit of the trace-free protocol is illustrated in Fig. 1b featuring one outer cycle and corresponding inner cycles. Additionally, as shown in Fig. 1b, two single-photon detectors (SPDs), denoted as $D_{0}$ and $D_{1}$, are positioned at the end of Alice’s outer interferometer cycles to detect bit 0 or 1 without photons traversing across the channel, while detector $D_{f}$ is placed at the end of the inner cycles to observe whether photons enter into the transmission channel.
The information transmitted is determined by Bob’s operation of switches (SWs), which act as obstructing objects. If Bob closes the switches to obstruct the evolution, choosing ${\hat{P}}={\hat{P}_{1}}$ in Fig. 1a, the quantum state of the photon is kept as initial and will be detected by $D_{0}$ for bit 0. Conversely, if Bob opens the switches to allow the photon to pass, choosing ${\hat{P}}={\hat{P}_{2}}$ in Fig. 1a, its quantum state evolves through interference and will be detected by $D_{1}$ for bit 1. As the numbers of $M$ and $N$ increase, the probability of successful transmission approaches 100% in theory due to the quantum Zeno effect, as shown in Fig. 1c.
@@ II.2 Experimental implementation
Although the probability of detecting the correct bit approaches 100% under infinite cycles of the outer and inner interferometers, it is impractical to realize extremely large numbers of $M$ and $N$ in an experiment. The core of the experimental design lies in selecting feasible values for $M$ and $N$, which determine the overall structure and success probabilities. In the modified protocol for counterfactual communication, a larger value of $M$ results in a lower error probability for transmitting bit 0, as given by $1-cos^{2}(\pi/2M)$. Regarding bit 1, the probability of transmission error depends on both values of $M$ and $N$. Given that larger values of $M$ and $N$ lead to a more complex structure with reduced interference visibility and increased channel loss, it is essential to identify appropriate values for $M$ and $N$ to maintain relatively high success probabilities while ensuring stable interference with acceptable channel loss. The success probabilities for transmitting bits 0 and 1 are simulated according to the modified protocol for counterfactual communication, as shown in Fig. 1c, where $S_{0}$ represents success probability for bit 0 and $S_{1}$ represents those for bit 1.
@@
In addition to the systematic scheme and on-chip structure mentioned above, another critical factor for achieving the protocol is the precise control of phases applied by all the PSs on the chip. To accomplish this, each PS is pre-calibrated to establish the relationship between the phase angle and the applied voltage, based on the thermal-optical effect. Furthermore, due to variations in the number of multi-mode interferometers (MMIs) in the two arms of tandem interferometers, phase compensation and loss offset are implemented for these arms (further details in Section IV). This optimization of phase implementation ensures accuracy in control and precision in success probability of information transmission for counterfactual communication. Consequently, the interference visibility within the entire structure, defined as the ratio of the interferometer’s minimum output intensity to its maximum output intensity, given by $V=\frac{I_{max}-I_{min}}{I_{max}+I_{min}}$, is maintained at over 99% for hours.
@@ III Discussion
Specifically, in the experiment, the SNSPDs, $D_{f1}$ and $D_{f2}$, are placed at the end of each inner cycle as Fig. 2(a) shown to check whether the photons entering the transmission channel in practice. We consider the click of $D_{f1}$$D_{f2}$ while $D_{0}$ ($D_{1}$) clicks simultaneously as a violation of counterfactual property. The coincident counts of $D_{f1}$$D_{f2}$ and $D_{0}$ ($D_{1}$) are summed up as the detection rate of the $D_{f}$ for bit 0 (bit 1). The experimental results illustrates that the detection rate of $D_{f}$ is $0.17\%$ in average ($0.09\%$ for bit 0 and $0.25\%$ for bit 1 respectively) while the transmission rate achieved in our experiment is $74.2\%$ for logic 0 and $85.1\%$ for logic 1. In our implementation of counterfactual communication, the high rate of $D_{0}$$D_{1}$ while low detection of $D_{f}$s is achieved by the long cycle interferometers, which matches well with the theoretical protocol.
@@
Via the quantum Zeno effect, our work implements the modified protocol with a long interference chain of outer and inner cycles via quantum nanophotonic chip, which guarantees the interference visibility and high success probabilities for bit 0 and bit 1. Benefiting from integrated circuit manufacturing, packaging technology, and booming quantum detection system, we adopt the integrated scheme to accomplish this experiment with higher ($99\%$) and lower error rate, compared with those implementations in free space [17, 21]. Moreover, to practically verify the counterfactual property, $D_{f}$ is appended at the end of the inner cycles to monitor photons’ leakage. Thus, the modified weak-trace-free counterfactual communication based on the quantum Zeno effect is fully realized. Furthermore, compared with the other implementation of modified counterfactual communication protocol published recently [21], our work implements the quantum Zeno effect, reduces the transmission time cost, and accelerates the transmission speed.
@@ IV Method
Figure 3: The whole experiment platform. Before entering the chip, pump light is produced by the laser and is adjusted with several steps, including spectral filtering by WDM and polarization modification by PC. The parameters of interferometer on chip are adjust by the signal DAC with Alice and Bob. Finally, the output photons are adjusted again and are detected by the SNSPD and counted with the TDC.
@@ IV.2 Phase compensation and loss offset
In this implementation, the on-chip components bring practical imperfections, deviating from the originally designed performance. Thus, phase compensation and loss offset are applied to optimize the interference. To realize the BSs controlling the outer and inner cycles’ interference, varied numbers of MMIs are placed in the four paths of the optical scheme. Since the MMI present a different reflectivity from the waveguide, the channel losses of the four paths are not identical due to the difference in the number of MMIs. In order to overcome the problem, firstly, we measure the loss of each path and calculate the loss shifting among the four paths. Then, the reflectivity of BS is altered according to the loss shifting, offsetting the different channel losses. Thus, the total channel loss is optimized to 11 $1\text{\,}\mathrm{dB}$ for bit 0 and $17\text{\,}\mathrm{dB}$ for bit 1. Because of the different schemes of interference for detecting bit 0 and bit 1, the channel loss varies in these two cases.
Moreover, due to using the phase shifters based on thermal-optical effect, heat crosstalk makes a difference on PSs’ modulation. The parameters when all the PSs are heated have a gap to those in calibration separately. In addition, the photons in path 2 have a phase shifting after $BS_{N}$s and more MMIs, compared to that in path 1. An adjustable PS in path 2 is used to modify the phase shifting for the outer cycles, which is calibrated under the condition that other modulators are set as working parameters. Consequently, the interference visibility is maintained over $99\%$ for 5 hours.
@@ Acknowledgements
This work was supported by National Natural Science Foundation of China (62371459, 62061136011, 62105366 and 62405368), the Research Fund Program of the State Key Laboratory of High Performance Computing (Grant No.202201-02). X.Q. acknowledges support from the National Natural Science Foundation of China (Grant No.62075243), the National Natural Science Foundation of China “Mathematical Basic Theory of Quantum Computing” special project (Grant No.12341103), and the National Young Talents Program.

### v2

@@ Abstract
keywords: counterfactual communication, quantum Zeno effect, quantum photonic integration
@@ I Introduction
The high-efficiency IFM is employed in various applications, such as quantum interrogation [17], quantum tomography [25, 12, 11], quantum computation [23, 15, 24, 26], and quantum cryptography [29, 18]. Importantly, this physical phenomenon is highly suitable for secure communication known as counterfactual communication, wherein information is conveyed without the transmission of any particles through the channel. As shown in Fig. 1(a), in classical communication, the communicating parties (Alice and Bob) must transmit information via physical particles passing through the channel, which may be eavesdropped by Eve. In counterfactual communication, however, Alice emits a quantum state to herself, and Bob conveys information by altering the evolution outcome of this state via $\hat{P}$. Throughout the entire process, no particles pass through the transmission channel. Thus the eavesdropper Eve cannot intercept any particles from the channel, and consequently cannot obtain the information. Furthermore, by utilizing the quantum Zeno effect, the success rate of counterfactual communication can be increased to 100% in theory [27].
@@ II.1 Counterfactual communication protocol
Figure 1: The schematic and structure of the upgraded counterfactual communication protocol. (a) During the classical communication process, Eve can obtain the information from Alice to Bob by intercepting the information-carried particles. However, since Alice and Bob do not transmit any particles during the counterfactual communication, Eve can not acquire any information. (b) The updated protocol for the counterfactual communication. (c) The success probabilities of transmitting bit 0 and bit 1 are simulated with various $M$ and $N$. Balancing the trade-off between feasibility of implementation and success probabilities of transmission, we choose $M=3$ and $N=6$ in our experimental scheme.
@@
Therefore, the specific optical scheme and the operation of the upgraded protocol are as follows, shown in Fig. 1(b). The outer cycles of interferometers consist of $M$ beam splitters ($BS_{M}$s) with the reflectivity of $cos^{2}(\pi/2M)$ each. Serving as one arm of the outer interferometer, the inner cycles contain two sets of tandem interferometers connected by a two-sided mirror. Each set of inner interferometers comprises $N$ beam splitters ($BS_{N}$s) with a reflectivity of $cos^{2}(\pi/2N)$ for each $BS_{N}$. In accordance with the quantum Zeno effect, the photon state remains stable within the nested cycles of interferometers. The structural unit of the trace-free protocol is illustrated in Fig. 1(b) featuring one outer cycle and corresponding inner cycles. Additionally, as shown in Fig. 1(b), two single-photon detectors (SPDs), denoted as $D_{0}$ and $D_{1}$, are positioned at the end of Alice’s outer interferometer cycles to detect bit 0 or 1 without photons traversing across the channel, while detector $D_{f}$ is placed at the end of the inner cycles to observe whether photons enter into the transmission channel.
The information transmitted is determined by Bob’s operation of switches (SWs), which act as obstructing objects. If Bob closes the switches to obstruct the evolution, choosing ${\hat{P}}={\hat{P}_{1}}$ in Fig. 1(a), the quantum state of the photon is kept as initial and will be detected by $D_{0}$ for bit 0. Conversely, if Bob opens the switches to allow the photon to pass, choosing ${\hat{P}}={\hat{P}_{2}}$ in Fig. 1(a), its quantum state evolves through interference and will be detected by $D_{1}$ for bit 1. As the numbers of $M$ and $N$ increase, the probability of successful transmission approaches 100% in theory due to the quantum Zeno effect, as shown in Fig. 1(c).
@@ II.2 Experimental implementation
Although the probability of detecting the correct bit approaches 100% under infinite cycles of the outer and inner interferometers, it is impractical to realize extremely large numbers of $M$ and $N$ in an experiment. The core of the experimental design lies in selecting feasible values for $M$ and $N$, which determine the overall structure and success probabilities. In the modified protocol for counterfactual communication, a larger value of $M$ results in a lower error probability for transmitting bit 0, as given by $1-cos^{2}(\pi/2M)$. Regarding bit 1, the probability of transmission error depends on both values of $M$ and $N$. Given that larger values of $M$ and $N$ lead to a more complex structure with reduced interference visibility and increased channel loss, it is essential to identify appropriate values for $M$ and $N$ to maintain relatively high success probabilities while ensuring stable interference with acceptable channel loss. The success probabilities for transmitting bits 0 and 1 are simulated according to the modified protocol for counterfactual communication, as shown in Fig. 1(c), where $S_{0}$ represents success probability for bit 0 and $S_{1}$ represents those for bit 1.
@@
In addition to the systematic scheme and on-chip structure mentioned above, another critical factor for achieving the protocol is the precise control of phases applied by all the PSs on the chip. To accomplish this, each PS is pre-calibrated to establish the relationship between the phase angle and the applied voltage, based on the thermal-optical effect. After these calibrations, we characterize the interference visibility of 27 MZIs in the cascade. The results show visibilities ranging from 99.42% to 99.99%, with a mean of 99.88% and a standard deviation of 0.0013, demonstrating excellent and consistent performance across all interferometers. Furthermore, due to variations in the number of multi-mode interferometers (MMIs) in the two arms of tandem interferometers, phase compensation and loss offset are implemented for these arms (further details in Sec. IV). This optimization of phase implementation ensures accuracy in control and precision in the success probability of information transmission for counterfactual communication. After global loss and phase compensation, the overall interference visibility of the entire structure is measured at a representative operating point to be 99.67%, confirming that the system achieves near-unity visibility under optimized conditions.
@@ III Discussion
Specifically, in the experiment, the SNSPDs, $D_{f1}$ and $D_{f2}$, are placed at the end of each inner cycle as Fig. 2(a) shown to check whether the photons entering the transmission channel in practice. We consider the click of $D_{f1}$$D_{f2}$ while $D_{0}$ ($D_{1}$) clicks simultaneously as a violation of counterfactual property. The coincident counts of $D_{f1}$ / $D_{f2}$ and $D_{0}$ ($D_{1}$) are summed as the detection rate of $D_{f}$ for bit 0 (bit 1). This coincidence counts utilizes raw detection data without subtracting dark counts. This ensures the reported $D_{f}$ detection rate remains a conservative upper bound, explicitly accounting for potential background noise within the results. The experimental results illustrate that the detection rate of $D_{f}$ is $0.17\%$ in average ($0.09\%$ for bit 0 and $0.25\%$ for bit 1 respectively) while the transmission rate achieved in our experiment is $74.2\%$ for logic 0 and $85.1\%$ for logic 1.
[…]
@@
Via the quantum Zeno effect, our work implements the modified protocol with a long interference chain of outer and inner cycles via quantum nanophotonic chip, which guarantees the interference visibility and high success probabilities for bit 0 and bit 1. Benefiting from integrated circuit manufacturing, packaging technology, and booming quantum detection system, we adopt the integrated scheme to accomplish this experiment with lower error rate, compared with those implementations in free space [7, 21]. Moreover, to practically verify the counterfactual property, $D_{f}$ is appended at the end of the inner cycles to monitor photons’ leakage. Thus, the modified weak-trace-free counterfactual communication based on the quantum Zeno effect is fully realized. Furthermore, compared with the other implementation of modified counterfactual communication protocol published recently [21], our work implements the quantum Zeno effect, reduces the transmission time cost, and accelerates the transmission speed.
For future work, it is worth considering the practical scalability limits of the nested interferometer architecture as $M$ and $N$ increase, along with potential strategies to mitigate these challenges. First, with $M$ and $N$ increasing, leading to a larger chip footprint and a reduced success yield [26], which is addressable through compact component designs, advanced wafer-scale manufacturing, or employing distributed chip architectures. Second, thermal crosstalk between adjacent thermo-optic PSs can be reduced using optimized bias calibration algorithms [6, 19], cascaded sub-array designs, or electro-optic modulation techniques. Third, control complexity is manageable through intelligent control algorithms and time-division multiplexing architectures [19]. These approaches collectively offer pathways toward scaling this architecture for future large-scale implementations.
@@ IV Method
Figure 3: The whole experiment platform. Before entering the chip, pump light is produced by the laser and is adjusted with several steps, including spectral filtering by WDM and polarization modification by PC. The parameters of the on-chip interferometer are adjusted by the signal DAC with Alice and Bob. Finally, the output photons are adjusted again and are detected by the SNSPD and counted with the TDC.
@@ IV.2 Phase compensation and loss offset
Due to the difference in reflectivity between MMIs and waveguides, as well as the varying number of MMIs in different paths, the transmission losses of different paths are not identical. This difference in loss would be accumulated with the increased number of M and N, which limits the scalability of the scheme if no compensation is conducted. To address this issue, we first quantify path imbalances by injecting classical light and measure the output optical power for the bit 0 and bit 1 paths, identifying transmission losses of 11 $1\text{\,}\mathrm{dB}$ and $17\text{\,}\mathrm{dB}$, respectively. Based on these measurements, we adjust the MZI voltage to tune the effective reflectivity of the BSs, thereby compensating for the loss discrepancy. Although inherent losses vary due to differing path lengths and component counts, the pre-adjusted MZI rebalances the photon’s transmitting probabilities to ensure high-visibility interference.
PSs are based on the thermo-optical effect and introduce additional phase shifts compared to waveguides. This leads to thermal crosstalk and phase shifts, which affect the interference effect. Moreover, compared to the outer loop path, the inner loop path experiences more severe phase shifts after passing through the $BS_{N}$ and more MMIs. Therefore, an adjustable PS is used in the inner loop path to correct the phase shift of the outer loop path, and this calibration is performed under the condition that other modulators are set to working parameters. This optimization of phase implementation ensures the accuracy of control and the precision of the success probability of information transmission in counterfactual communication. As a result, the interference visibility within the entire structure reaches 99.67%, defined as the ratio of the interferometer’s minimum output intensity to its maximum output intensity, given by $V=\frac{I_{max}-I_{min}}{I_{max}+I_{min}}$. The long-term stability of the system supports the sustained high success probabilities (74.2% for bit 0 and 85.1% for bit 1) over a 5-hour continuous experiment.
@@ Acknowledgements
This work was supported by National Natural Science Foundation of China (62371459, 62061136011, 62105366 and 62405368), the Quantum Science and Technology—National Science and Technology Major Project (2021ZD0300704). X.Q. acknowledges support from the National Natural Science Foundation of China (Grant No.62075243), the National Natural Science Foundation of China “Mathematical Basic Theory of Quantum Computing” special project (Grant No.12341103), and the National Young Talents Program.